# Jaike van Twiller1    Yossiri Adulyasak2    Erick Delage2    Djordje Grbic1    Rune Møller Jensen1

### v3

@@ 4.1 Formal MDP
4.2 Decomposed MDP
@@
The formal MDP defines an effective action space of size $|{X}|\propto|B|\cdot|D|\cdot|K|\cdot|P|$, where each action ${x}_{p}$ determines how cargo types and transport options are placed on the vessel per port. However, this action space is large, which can hinder learning efficiency Kanervisto et al. (2020).
@@
To address this, we decompose the formal MDP into granular, sequential steps based on an index $(i,j,k)\;\forall(i,j)\in\textit{TR},k\in K$. Instead of placing all transport and cargo types simultaneously, we take a decomposed action for all transports $(p,j)$ and cargo types $k$ at port $p$, then departing to a new port. This reduces the action space to $|X|=|B|\cdot|D|$, while unfolding transports and cargo types over an extended time horizon $t\in H=\{0,1,\dots,T_{\textit{seq}}\}$ with $T_{\textit{seq}}=|K|\cdot|\textit{TR}|.$
@@
State. The state $s_{t}=(u_{t},q_{t},\zeta)$ depends on time step $t$, where $u_{t}\in\mathbb{R}^{n_{u}}_{\geq 0}$ is vessel utilization, and $q_{t}\in\mathbb{R}^{n_{q}}_{\geq 0}$ is realized demand. The environment parameter $\zeta$ remains unchanged. Given the time $t$, however, we can extract relevant parameters from $\zeta$, such as $(\textit{pol}_{t},\textit{pol}_{t},k_{t}),\textit{rev}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}$.
@@
Action. Action $x_{t}\in\mathbb{R}^{n_{c}}_{\geq 0}$ assigns real number of containers to utilization $u_{t}$ for step $t$. Each action is subject to $\textit{PH}(s_{t})=\{x_{t}\in\mathbb{R}^{n_{c}}_{>0}:A(s_{t})x_{t}\leq b(s_{t})\}$. Here, $A(s_{t})\in\mathbb{R}^{m_{c}\times{n_{c}}}$ is the constraint matrix, $b(s_{t})\in\mathbb{R}^{m_{c}}$ is the bound vector, and $m_{c}$ is the number of constraints. It is also worth noting that Constraints (3)-(8) are reformulated to fit feasible region $\textit{PH}(s_{t})$.
@@
Transition. At each time step $t$, the transition includes loading, where $x_{t}$ is added to $u_{t}$. However, discharging and demand realization occur only when arriving at a new port, indicated by $t\in T_{\text{new port}}$.
@@
Reward. The revenue at step $t$ is computed as $\textit{rev}(\textit{pol}_{t},\textit{pod}_{t},k_{t})\min(\mathbf{1}^{\top}x_{t},q_{t}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}).$ However, costs depend on knowing all loading operations at port $p$, which is aggregated in utilization $u_{t}$ at the last step of the port $t\in T_{\text{leave port}}$. As a result, the cost signal is sparse, being evaluated only once per port $p$, rather than at each step.
5 Proposed Architecture
Our approach consists of several components: an encoder-decoder model parameterized by $\theta$, an actor-critic DRL method, and a feasibility projection layer. The encoder-decoder model shown in Figure 2 employs a look-ahead policy $\pi_{\theta}(x|s_{t})$ conditioned on state $s_{t}$ and parameterized by mean $\mu_{\theta}(s_{t})$ and standard deviation $\sigma_{\theta}(s_{t})$. By training in the decomposed MDP and optionally projecting the policy, we iteratively generate actions $x_{t}$ to construct solutions.
@@ 5 Proposed Architecture
Figure 2: Deep Reinforcement Learning Architecture with Feasibility Projection for Actor-Critic Methods
5.1 Encoder-Decoder Model
Figure 3 presents our encoder-decoder model, based on that of Kool et al. (2019), with modifications highlighted below.
Figure 3: Layers of the encoder and the actor-critic decoder.
Cargo Embedding. The cargo embedding $e^{\text{cargo}}(\zeta)$ parameterized by $\theta_{\textit{cr}}$ maps cargo-related episode information in $\zeta$ into a feature representation for the encoder. To enhance positional awareness, we subsequently apply sinusoidal positional encoding Vaswani et al. (2017).
@@ 5.1 Encoder-Decoder Model
Encoding Layers. An attention encoder $f(e^{\text{cargo}}(\zeta))$ parameterized by ${\theta_{\textit{enc}}}$ maps embedding $e^{\text{cargo}}(\zeta)$ to latent variable $z$ using multi-head attention (MHA) to identify revalant features dynamically Vaswani et al. (2017). Then, we use a feed-forward network (FFN) with ReLU activation, layer normalization, residual connections, and dropout.
@@
Context Embedding. The context embedding $e^{\text{cont}}(u_{t},z)$, parameterized by ${\theta_{\textit{co}}}$, focuses on time $t$, extracting time-specific features from the utilization $u_{t}$ and latent variable $z$. These features provide the MHA query with a representation of the vessel’s current condition, enabling the policy to make decisions based on the present context.
@@
Dynamic Embedding. The dynamic embedding $e^{\text{dynam}}(q_{t},z)$, parameterized by ${\theta_{\textit{dn}}}$, extracts features from demand $q_{t}$ in all steps of horizon $H$, capturing patterns across the entire episode. It combines real demand and latent information to produce keys and values for the MHA layer. As such, it accounts for global temporal trends, allowing it to anticipate future conditions and make long-term decisions.
@@
Actor Layers. As proposed in Kool et al. (2019), our actor decoder $g(e^{\text{cont}}(u_{t},z),e^{\text{dynam}}(q_{t},z))$ is an attention model parameterized by $\theta_{\textit{act}}$. However, our actor takes input from the context and dynamic embedding. An MHA layer with a forward-looking mask bases decisions on steps ${t,t+1,\dots,T_{\textit{seq}}}$ to anticipate future events. Subsequently, an FFN extracts features, after which a pointer mechanism performs a soft selection of relevant steps using key-value inputs Vinyals et al. (2015). A softplus activation outputs positive logits $({\mu}_{\theta}(s_{t}),{\sigma}_{\theta}(s_{t}))$.
@@
Critic Layers. Our critic model $V(s_{t},z)$, parameterized by $\theta_{\textit{cri}}$, estimates the value of state $s_{t}$ and latent variable $z$ through an FFN outputting $V_{\theta_{\textit{cri}}}(s_{t},z)\in\mathbb{R}$.
@@
Action Policy. The actor logits parameterize a stochastic policy $\pi_{\theta}(x|s_{t})=\mathcal{N}(({\mu}_{\theta}(s_{t}),{\sigma}_{\theta}(s_{t}))$, which allows action sampling $x_{t}\sim\pi_{\theta}(x|s_{t})$ to construct solutions.
5.2 Feasibility Regularization in Actor-Critic Loss
Standard actor-critic methods can efficiently and stably learn policies that require implicit feasibility in the action space Schulman et al. (2017); Haarnoja et al. (2018), while explicit constraints are often integrated by feasibility regularization (FR) in the actor loss Ding et al. (2020); Calvo-Fullana et al. (2021). Equation (11) defines a composite loss that combines the actor loss $\mathcal{L}_{\text{actor}}(\theta)$ with a soft regularization term $\mathcal{L}_{\text{feas}}(\theta)$. FR is defined in Equation (11), where $\lambda_{f}$ controls the regularization strength, and policy samples $x_{\theta}(s_{t})$ are assumed to allow gradient flow to update parameters $\theta$, as is the case in SAC Haarnoja et al. (2018). For algorithms that do not include actions in the computation graph, e.g., PPO Schulman et al. (2017), we substitute $x_{\theta}(s_{t})$ with $\mu_{\theta}(s_{t})$.
Determining Lagrangian multipliers ($\lambda_{f}$) for multiple constraints with varying scales is already challenging for static feasible regions. It requires costly hyperparameter tuning to balance a trade-off between objectives and feasibility effectively. As a result, FR is a naive approach when applied to dynamic, state-dependent feasible regions $\textit{PH}(s_{t})$ Calvo-Fullana et al. (2021); Donti et al. (2021). However, FR remains useful as a comparative baseline.
@@ 5.2 Feasibility Regularization in Actor-Critic Loss
$\mathcal{L}^{\text{feas}}(\theta)=$ $\;\mathbb{E}_{t}\left[\left(\max(A(s_{t}){x}_{\theta}(s_{t})-b(s_{t}))_{>0}\right)^{2}\right]$ (11)
5.3 Feasibility Layers
To extend beyond FR, we leverage differentiable projection layers for constraint minimization. We incorporate two constraint-specific and one general layer(s). Note that a single action may be infeasible while combining such actions can achieve feasibility. Therefore, we prioritize violation minimization over strict enforcement to allow for flexibility.
Furthermore, non-linear transformations of continuous distribution samples change their probability density. We account for this transformation using Jacobian adjustments to the policy’s log probabilities Bishop (2006).
Technical details on feasibility layers are provided in Appendix B of supplementary material.
Weighted Scaling Projection. Function (12) defines the weighted scaling layer (WS), which normalizes vector $x$ if the sum of $x$ exceeds scalar $y$. This preserves relative proportions of elements in $x$ while enforcing $x$ to sum to scalar $y$.
Policy Clipping. We can apply function $\mathcal{C}(x,\textit{lb}_{\textit{pc}},\textit{ub}_{\textit{pc}})=\max(\min(x,\textit{ub}_{\textit{pc}}),\textit{lb}_{\textit{pc}})$ to enforce element-wise limits on vector $x$, thereby performing policy clipping (PC) on actions. However, PC is only applicable to box constraints.
Violation Projection. In Algorithm 1, we define a violation projection (VP) layer that reduces inequality constraint violations by shifting point $x$ closer to the feasible region of some convex polyhedron $\textit{PH}=\{x\in\mathbb{R}^{n}_{>0}:Ax\leq b\}$, where $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$. The inequality $Ax\leq b$ ensures PH is convex, enabling gradient-based reduction of violation $\mathcal{V}(x)$ Boyd and Vandenberghe (2004). To measure feasibility, we compute the element-wise violation $\mathcal{V}(x)=(Ax-b)_{>0}$, where each $\mathcal{V}(x)_{m_{i}}>0$ indicates the violation of constraint ${}_{m_{i}}$, and $\mathcal{V}(x)_{m_{i}}=0$ indicates satisfaction of that constraint. To minimize violations, we iteratively update $x$ by applying gradient descent to reduce the violation term $\|\mathcal{V}(x)\|_{2}^{2}$, representing the squared distance from $x$ to PH. Equation (13) updates $x$ with a step size $\eta_{v}$, which is rewritten to the update function found in Algorithm 1.
@@ 5.3 Feasibility Layers
During training, the VP layer continues for a fixed number of epochs. However, we include a stopping criteria during inference, which continues until the change in total violation $\mathbf{1}^{\top}\mathcal{V}(x^{\prime})-\mathbf{1}^{\top}\mathcal{V}(x)$ is below a threshold $\delta_{v}$. As a result, we $x$’s distance to the feasible region and incorporate constraint awareness into initially unconstrained policies.
Algorithm 1 Violation Projection Layer
0:  $x\in\mathbb{R}^{n}_{>0}$, parameters $(A,b,\eta_{v},\delta_{v})$
@@
Methods Testing ($\boldsymbol{N=30}$) Generalization ($\boldsymbol{N=30}$)
Alg. Model F.M. Ob. (\$) Time (s) F.I. (%) F.O. (\$) Ob. (\$) Time (s) F.I.(%) F.O. (\$)
SAC AM FR 1113.03 12.63 0.00 1065.80 1211.87 15.35 0.00 1041.42
SAC AM-P FR/VP 1318.05 15.20 100.00 - 1288.68 14.18 100.00 -
SAC AM-P VP 1447.69 15.03 100.00 - 1411.34 13.33 100.00 -
SAC AM-P FR/WS/PC 1373.24 13.34 100.00 - 1340.12 14.42 100.00 -
SAC AM-P WS/PC 1494.22 13.12 100.00 - 1482.40 12.91 100.00 -
PPO AM FR 1842.46 11.74 0.00 1063.24 1830.95 12.60 0.00 1066.69
PPO AM-P FR/VP 1355.45 15.60 100.00 - 1321.90 14.91 100.00 -
PPO AM-P VP 1318.10 14.51 100.00 - 1282.10 14.64 100.00 -
PPO AM-P FR/WS/PC 1369.03 14.29 100.00 - 1330.10 14.17 100.00 -
PPO AM-P WS/PC 1471.40 13.75 100.00 - 1455.98 13.21 100.00 -
CPL SMIP-NA - 1053.02 8434.58 100.00 - 1023.04 8434.94 100.00 -
CPL SMIP-PI* - 1713.73 40.75 100.00 - 1680.51 38.93 100.00 -
Table 1: Experimental results comparing DRL methods (SAC/PPO) with a vanilla (AM) or projected attention model (AM-P) and feasibility mechanisms (F.M.): feasibility regularization (FR), weighted scaling (WS), policy clipping (PC) and violation projection (VP). We compare SMIP-NA (non-anticipation) as a baseline, and SMIP-PI (* assumes perfect information which is unrealistic) as an expected upper bound, both solved with CPLEX (CPL). Average performance metrics on $N$ instances include objective value in profit (Ob.), inference time in seconds (Time), percentage of feasible instances (F.I.), and objective value with feasibility recovery (F.O.). Note that † indicates infeasible objectives. Generalization performance is evaluated on unseen, out-of-distribution instances based on a uniform distribution.
6 Experimental Results
@@
We compare our policies against baselines on in-distribution and out-of-distribution instances, evaluating the objective value, computational cost and feasibility. Additionally, an ablation study analyzes the impact of feasibility mechanisms and gradient flow from actions, while managerial insights discuss the value of information, computational cost, and adaptiveness to varying levels of uncertainty.
6.1 Experimental Setup
@@
Instance Generation. Training instances are sampled from a Gaussian distribution $\mathcal{N}(\mu^{(i,j,k)},\sigma^{(i,j,k)})\;\forall(i,j)\in\textit{TR},k\in K$, where $\mu^{(i,j,k)}$ is randomly generated and $\sigma^{(i,j,k)}=\textit{CV}\mu^{(i,j,k)}$ controls spread via the coefficient of variation CV. A continuous uniform distribution generates out-of-distribution instances $\mathcal{U}(\textit{lb}^{(i,j,k)},\textit{ub}^{(i,j,k)})\;\forall(i,j)\in\textit{TR},k\in K$ to test generalization. Details are in Appendix C.
Feasibility Implementation. We evaluate feasibility by integrating multiple mechanisms to address all constraints in $\textit{PH}(s_{t})$. Starting with a baseline using FR, we introduce two alternatives: a constraint-specific (FR/WS/PC) and a general projection approach (FR/VP). We then remove FR from both to assess its impact. WS and PC are not analyzed separately, as it leaves many constraints unaddressed.
SMIP Baseline. We compare our approach against two common techniques in stochastic programming: SMIP-NA (stochastic MIP without anticipation) and SMIP-PI (stochastic MIP with perfect foresight, serving as an upper bound). The scenario trees and MIP models are shown in Appendix D.
Runs. Training occurs offline on simulated instances for a 1,000 TEU vessel over a 4-port voyage. GPU-based experiments use an NVIDIA RTX A6000, and CPU-based runs use an AMD EPYC 9454 48-core processor. Implementation details and additional experiments are provided in Appendix E.
6.2 Policy Performance
Table 1 compares the performance of different solution approaches. Our projected attention models (AM-P) outperform SMIP-NA, achieving around 30-35% higher objective values and 500× faster computations. The AM with FR misleadingly attains higher profits than SMIP-NA due to infeasibility. If we recover feasibility, then profits are significantly reduced to the level of SMIP-NA. In contrast, each AM-P finds feasible solutions with a higher profit than AM with FR, highlighting FR’s limitations. All AM-P policies generalize well, with SAC and PPO showing an average profit reduction of 2% and 3% on unseen uniform instances, similar to SMIP-NA and SMIP-PI. These results suggest that AM-P can leverage imperfect information to anticipate future uncertainty and ensure feasibility by projection layers beyond FR.
Ablation Study. Table 1 analyzes the impact of feasibility mechanisms. Removing FR from FR/VP increases profit for SAC and slightly decreases profit for PPO, without diminishing feasibility. When FR is removed from FR/WS/PC, profit increases, and both DRL approaches continue to maintain feasibility. This outcome is surprising, given the absence of a direct mechanism to account for stability. However, from our domain understanding, we know container vessels generally become more stable as utilization increases in all locations. Replacing SAC with PPO, which eliminates action gradient flow, reduces feasibility in FR, though projection layers mitigate this effect. These findings highlight the impact of feasibility mechanisms on the objective value and feasibility.
6.3 Managerial Insights

### v5

@@ Introduction
Figure 2: Illustration of the decomposed MDP for MPP. At port $p$, the vessel discharges cargo, observes realized demand $q_{t}$, then sequentially loads container type–POD pairs before sailing to the next port p+1p{+}1. Colors indicate PODs.
Decomposed Markov Decision Process
@@
The MDP has an action space of size $|{X}|\propto|B|\cdot|D|\cdot|K|\cdot|P|$, where each action ${x}_{p}$ determines how cargo types and transport options are placed on the vessel per port. However, this action space is large, which can hinder learning efficiency (Kanervisto, Scheller, and Hautamäki 2020).
@@
To address this, we decompose the MDP into granular, sequential steps based on an index $(i,j,k)\;\forall(i,j)\in\textit{TR},k\in K$. Figure 2 illustrates the decomposed MDP for a single port. Instead of placing all transport and cargo types simultaneously, we take a decomposed action for all transports $(p,j)$ and cargo types $k$ at port $p$, then departing to a new port. This reduces the action space to $|X|=|B|\cdot|D|$, while unfolding transports and cargo types over an extended time horizon $t\in H=\{0,1,\dots,T_{\textit{seq}}\}$ with $T_{\textit{seq}}=|K|\cdot|\textit{TR}|.$
State.
@@
The state $s_{t}=(u_{t},q_{t},\zeta)$ depends on time step $t$, where $u_{t}\in\mathbb{R}^{n_{u}}_{\geq 0}$ is vessel utilization, and $q_{t}\in\mathbb{R}^{n_{q}}_{\geq 0}$ is realized demand. The instance $\zeta$ remains unchanged. Given the time $t$, however, we can extract relevant parameters from $\zeta$, such as $(\textit{pol}_{t},\textit{pol}_{t},k_{t}),\textit{rev}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}$.
Action.
@@
Action $x_{t}\in\mathbb{R}^{n_{c}}_{\geq 0}$ assigns real number of containers to utilization $u_{t}$ for step $t$. Each action is subject to $\textit{PH}(s_{t})=\{x_{t}\in\mathbb{R}^{n_{c}}_{>0}:A(s_{t})x_{t}\leq b(s_{t})\}$. Here, $A(s_{t})\in\mathbb{R}^{m_{c}\times{n_{c}}}$ is the constraint matrix, $b(s_{t})\in\mathbb{R}^{m_{c}}$ is the bound vector, and $m_{c}$ is the number of constraints. It is also worth noting that Constraints (3)-(8) are reformulated to fit feasible region $\textit{PH}(s_{t})$.
Transition.
@@
At each time $t$, the transition includes loading, where $x_{t}$ is added to $u_{t}$. Discharges and demand realizations occur when arriving at a new port, indicated by $t\in T_{\text{new port}}$.
Reward.
@@
The revenue at step $t$ is computed as $\textit{rev}(\textit{pol}_{t},\textit{pod}_{t},k_{t})\min(\mathbf{1}^{\top}x_{t},q_{t}^{(\textit{pol}_{t},\textit{pod}_{t},k_{t})}).$ However, costs depend on knowing all loading operations at port $p$, which is aggregated in utilization $u_{t}$ at the last step of the port $t\in T_{\text{leave port}}$. As a result, the cost signal is sparse, being evaluated only once per port $p$, rather than at each step.
Proposed Architecture
Figure 3 shows our projected encoder-decoder model, which embeds instance, context, and dynamic data to obtain a stochastic look-ahead policy $\pi_{\theta}(x|s_{t})$ parameterized by mean $\mu_{\theta}(s_{t})$ and standard deviation $\sigma_{\theta}(s_{t})$. The policy output is projected by layer $\mathcal{P}$ to ensure feasibility, and adjusted actions are used in the decomposed MDP. We train the projected policy with actor-critic RL, learning to interact in the decomposed MDP with a critic evaluating actions.
@@
Figure 3: Deep reinforcement learning architecture with feasibility projection for actor-critic methods.
Encoder-Decoder Model
Our model, adapted from Kool, van Hoof, and Welling (2019) and extended with a critic, is illustrated in Figure 4.
Figure 4: Layers of the encoder and the actor-critic decoder
Instance Embedding.
The embedding $e^{\text{inst}}(\zeta)$, parameterized by $\theta_{\textit{cr}}$, maps problem instance $\zeta$ to a feature representation used by the encoder to condition policies on instances. Sinusoidal positional encoding is then applied to inform the model of element order in sequences (Vaswani et al. 2017).
Encoding Layers.
@@
An attention encoder $f(e^{\text{inst}}(\zeta))$ parameterized by $\theta_{\textit{enc}}$} maps embedding $e^{\text{inst}}(\zeta)$ to latent variable $z$ using multi-head attention (MHA) to identify relevant features dynamically and handle variable input size (Vaswani et al. 2017). Then, we use a feed-forward network (FFN) with ReLU activation, layer normalization, residual connections, and dropout.
Context Embedding.
@@
The embedding $e^{\text{cont}}(u_{t},z)$, parameterized by $\theta_{\textit{co}}$, extracts features from the current solution $u_{t}$ and the latent variable $z$. This embedding provides the MHA query, helping the attention mechanism focus on relevant information, with the solution context at time $t$, enabling the policy to make informed decisions on the present state.
Dynamic Embedding.
@@
The embedding $e^{\text{dyn}}(q_{t},z)$, parameterized by $\theta_{\textit{dn}}$, extracts features from demand $q_{t}$ across all time steps in the horizon $H$, capturing temporal patterns over the episode. By combining actual demand with latent information, it enables the attention mechanism to select (key) and retrieve (value) demand-related information over time for anticipating future conditions and long-term decision-making.
Actor Layers.
@@
Following (Kool, van Hoof, and Welling 2019), the actor decoder $g(e^{\text{cont}}(u_{t},z),e^{\text{dyn}}(q_{t},z))$ is an attention model (AM) parameterized by $\theta_{\textit{act}}$. The MHA layer uses context embedding as queries and dynamic embedding as keys and values encoding demand across the horizon, with a forward-looking mask on steps $\{t,t\!+\!1,\dots,T_{\textit{seq}}\}$ to link present context with future trends. An FFN extracts features, a pointer mechanism selects relevant future steps (Vinyals, Fortunato, and Jaitly 2015), and softplus activation then outputs positive logits $({\mu}_{\theta}(s_{t}),{\sigma}_{\theta}(s_{t}))$.
Critic Layers.
@@
Our critic model $V(s_{t},z)$, parameterized by $\theta_{\textit{cri}}$, estimates the value of state $s_{t}$ and latent variable $z$ through an FFN outputting $V_{\theta_{\textit{cri}}}(s_{t},z)\in\mathbb{R}$.
Action Policy.
@@
The actor logits parameterize a stochastic Gaussian policy $\pi_{\theta}(x|s_{t})=\mathcal{N}(({\mu}_{\theta}(s_{t}),{\sigma}_{\theta}(s_{t}))$, which allows action sampling $x_{t}\sim\pi_{\theta}(x|s_{t})$ to construct solutions.
Feasibility Regularization in Actor-Critic Loss
Actor-critic methods such as proximal policy optimization (PPO)  (Schulman et al. 2017) and soft actor critic (SAC)  (Haarnoja et al. 2018) are widely used in unconstrained problems with implicitly encoded feasibility. A common approach to handle explicit action constraints is to augment the actor loss with a soft penalty term, known as feasibility regularization (FR) (Ding et al. 2020; Calvo-Fullana et al. 2021).
The actor loss is then defined as:
@@
$\mathcal{L}_{\text{feas}}(\theta)$ $=\mathbb{E}_{t}\left[\left(A(s_{t})x_{\theta}(s_{t})-b(s_{t})\right)_{>0}\right]$ (11)
where $(\cdot)_{>0}$ denotes the ReLU function applied elementwise, penalizing only constraint violations, and $\lambda_{f}$ controls the strength of regularization. For algorithms that do not propagate gradients through the action (e.g., PPO), we substitute $x_{\theta}(s_{t})$ with the policy mean $\mu_{\theta}(s_{t})$.
However, tuning Lagrangian multipliers ($\lambda_{f}$) is challenging, especially with constraints of varying scales, even in static feasible regions. The problem worsens in dynamic, state-dependent regions $\textit{PH}(s_{t})$ (Calvo-Fullana et al. 2021; Donti, Rolnick, and Kolter 2021), where the balance between reward and feasibility shifts over time. Thus, FR is often limited in such settings, though it remains a useful baseline due to its simplicity and broad adoption.
Feasibility Projection Layers
To improve over FR, we incorporate differentiable projection layers that project sampled actions $x_{t}$ before use in the decomposed MDP and policy update. These layers guarantee feasible actions and maintain gradient flow by correcting non-linear density changes via Jacobian adjustments to policy log-probabilities (Bishop 2006). Existing differentiable convex optimization layers (Agrawal et al. 2019) use implicit maps with intractable Jacobians, leading to biased gradients that hurt learning accuracy and high computational cost that limits efficiency. In contrast, we design unbiased projections with tractable Jacobians that enable corrections of log-probabilities, including one general convex layer and two constraint-specific variants. Details are in Appendix B.
Unbiased Violation Projection.
Algorithm 1 introduces the unbiased violation projection layer (UVP), which iteratively moves a sampled action $x$ closer to the feasible set defined by a convex polyhedron $\textit{PH}=\{x\in\mathbb{R}^{n}_{>0}:Ax\leq b\},$ where $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$. Rather than enforcing strict feasibility at each step, UVP minimizes total constraint violation using gradient descent on the violation norm $\|\mathcal{V}(x)\|_{2}^{2},\text{ where }\mathcal{V}(x)=(Ax-b)_{>0}$ which measures element-wise violations of constraints (Boyd and Vandenberghe 2004). Equation (12) provides the update rule with step size $\eta_{v}$, as implemented in Algorithm 1.
@@
As shown in Appendix B, UVP iteratively reduces constraint violations without requiring exact feasibility, converging to a point of minimal violation and improving feasibility likelihood. During training, it runs for a fixed number of epochs; at inference, it stops when the change in total violation falls below $\delta_{v}$. Although UVP alters actions, we show in Appendix B that it supports accurate learning by applying Jacobian-based log-probability corrections to account for the volume change induced by the transformation. This ensures distributional unbiasedness, i.e., the expected gradient after projection matches that before projection, which is crucial for correct and reliable learning with projection.
Algorithm 1 Unbiased violation projection layer
[…]
@@
Weighted Scaling Projection.
Function (13) defines the weighted scaling layer (WS), which normalizes vector $x$ if the sum of $x$ exceeds scalar $y$. This preserves relative proportions of elements in $x$ while enforcing $\mathbf{1}^{\top}x=y$.
Policy Clipping.
We can apply function $\mathcal{C}(x,\textit{lb}_{\textit{pc}},\textit{ub}_{\textit{pc}})=\max(\min(x,\textit{ub}_{\textit{pc}}),\textit{lb}_{\textit{pc}})$ to enforce element-wise limits on vector $x$, thereby performing policy clipping (PC) on actions. However, PC is only applicable to box constraints.
Experimental Results
@@
We compare across data distributions on objective value (Obj.), training and test time (Time), and percentage of feasible instances (Feas.). An ablation study analyzes feasibility methods and gradient flow from actions, while managerial insights examine information value, computational cost, adaptiveness to uncertainty, and scaling to longer voyages.
Experimental Setup
Instances.
@@
Training instances are Gaussian samples from $\mathcal{N}(\mu^{(i,j,k)},\sigma^{(i,j,k)})\;\forall(i,\!j)\in\textit{TR},k\in K$, with a randomized $\mu^{(i,j,k)}$ and spread $\sigma^{(i,j,k)}=\textit{CV}\mu^{(i,j,k)}$ controlled by coefficient of variation CV. Generalization instances are continuous uniform samples from $\mathcal{U}(\textit{lb}^{(i,j,k)},\textit{ub}^{(i,j,k)})\;\forall(i,j)\in\textit{TR},k\in K$. Details are found in Appendix C.
Feasibility Mechanisms.
Table 1 summarizes the feasibility mechanisms. Feasibility regularization (FR) uses a static composite loss, while primal-dual methods (PD) learn $\lambda_{f}$ for the same composite loss. AM with projection (AM-P) includes general (violation projection (UVP), convex program (CP) by Agrawal et al. (2019)), and constraint-specific layers (weighted scaling (WS), policy clipping (PC)). UVP, WS, and PC are evaluated standalone, but also with FR during training, and with CP-based inference.
Type Implementation Constraints
FR/PD Composite loss Entire $\textit{PH}(s_{t})$
UVP $\textit{UVP}(x_{t},\!A(s_{t}),\!b(s_{t}),\!\alpha_{v},\!\delta_{v})$ Entire $\textit{PH}(s_{t})$
CP $\textit{CP}(x_{t},\!A(s_{t}),\!b(s_{t}))$ (inf.) Entire $\textit{PH}(s_{t})$
WS $\mathcal{W}(x_{t},\!q_{t})$ Demand $q_{t}$
PC $\mathcal{C}(x_{t},\!0,\!c\!-\!u^{\prime}_{t}\textit{teu})$ Capacity $c$
Table 1: Feasibility mechanisms and relation to constraints
Runs and Baselines.
Offline training (4-port, 1000 TEU vessel) was conducted on an NVIDIA RTX A6000 GPU. Baselines span two categories: solving implicit stochastic environments (AM with FR or PD, AM-P with CP) and solving explicit scenario trees, (stochastic MIP without anticipation (SMIP-NA) and with perfect foresight (SMIP-PI) as an upper bound). SMIPs ran on an AMD EPYC 9454 48-core CPU. Implementation details are in Appendix D and E.
Methods Train Testing ($\boldsymbol{N=30}$) Generalization ($\boldsymbol{N=30}$)
Alg. Model F.M. Time (s) Obj. (\$) Time (s) Feas. (%) Obj. (\$) Time (s) Feas.(%)
SAC AM-P UVP 11,781 1,451.30 0.74 100.00 1,476.74 0.74 100.00
SAC AM-P FR/UVP 7,612 1,327.03 0.65 100.00 1,413.26 0.69 100.00
SAC AM-P UVP/CP 11,781 1,479.41 1.32 100.00 1,505.32 1.33 100.00
SAC AM-P WS/PC 9,015 1,474.17 0.67 100.00 1,500.57 0.67 100.00
SAC AM-P FR/WS/PC 7,201 1,370.14 0.68 100.00 1,399.58 0.68 100.00
SAC AM-P WS/PC/CP 9,015 1,479.02 1.32 100.00 1,504.69 1.32 100.00
SAC AM FR 2,443 1,188.90 0.76 0.00 1,196.70 1.12 0.00
SAC AM PD 5,205 1,077.84 0.76 0.00 1,085.58 0.87 0.00
SAC AM-P CP 85,031 1,482.56 11.74 100.00 1,509.05 11.80 100.00
PPO AM-P UVP 15,517 1,353.58 0.97 100.00 1,373.96 0.97 100.00
PPO AM-P FR/UVP 16,585 1,328.10 0.67 100.00 1,350.19 0.67 100.00
PPO AM-P UVP/CP 15,517 1,482.72 1.33 100.00 1,508.85 1.33 100.00
PPO AM-P WS/PC 21,860 1,453.53 0.67 100.00 1,471.66 0.67 100.00
PPO AM-P FR/WS/PC 4,885 1,331.02 0.67 100.00 1,353.19 0.67 100.00
PPO AM-P WS/PC/CP 21,860 1,480.45 4.70 86.67 1,508.78 4.65 93.33
PPO AM FR 5,800 1,294.31 0.75 0.00 1306.85 1.43 0.00
PPO AM PD 7,331 1,149.44 1.42 0.00 1,152.83 0.75 0.00
PPO AM-P CP 233,959 1,479.25 11.21 96.67 1,499.46 8.28 100.00
CPL SMIP-NA - - 1,053.02 8,434.58 100.00 1,023.04 8,434.94 100.00
CPL SMIP-PI* - - 1,713.73 40.75 100.00 1,680.51 38.93 100.00
Table 2: AM-P versus baselines with various algorithms (Alg.) and feasibility mechanisms (F.M.).
Policy Performance
Table 2 compares the performance of different methods. AM-P outperforms SMIP-NA with 30–40% higher objective profits and over 2 hours faster test time computation. AM with FR or PD shows inflated profits due to infeasible solutions; Appendix F confirms that feasibility recovery ensures profits of \$1,200 and lower. In contrast, AM-P mostly finds feasible solutions with higher profits, highlighting FR and PD’s limitations. Among feasibility methods, baseline CP yields high profit, but can be infeasible and requires much more training and test time. Except for PPO, UVP/CP and WS/PC/CP offer the best trade-off between profit, training, and test time, while UVP and WS/PC achieve lower profits with lower test time. Each AM-P generalizes well: SAC and PPO improve average profit by 3% and 2% on unseen test instances. Thus, our AM-P uniquely outperforms the state-of-the-art, demonstrating a general approach for stochastic, dynamic sequential problems with state-dependent constraints.
Ablation Study.
Table 2 also presents an ablation study. Removing FR from FR/UVP increases profit significantly for SAC and slightly for PPO without affecting feasibility. Similarly, removing FR from FR/WS/PC increases profit and maintains feasibility for both DRL methods. Though surprising without stability mechanisms, the performance aligns with the principle that vessels stabilize with higher utilization. Replacing SAC with PPO for action-gradient flow increases profit but also violations in FR and PD (see Appendix F), while projection mitigates this effect.
Managerial Insights